OpenStudy (anonymous):

The position vector for a particle moving on a helix is c(t)= (cost, sint, t^2). Find the speed of the particle at time t = 4pi

OpenStudy (anonymous):

what is the velocity vector?

OpenStudy (anonymous):

i have already plugged 4pi into each to get ( 1, 0, 16pi^2 ). just dont know what to do now

OpenStudy (turingtest):

given a position function, velocity is found by...?

OpenStudy (anonymous):

wouldn't velocity be c'(t)?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

that is another part or the question... haven't gotten there yet. was looking for a formula to find speed

OpenStudy (anonymous):

how is speed related to velocity?

OpenStudy (anonymous):

the anti-deriv. of velocity?

OpenStudy (turingtest):

position is the anti-derivative of velocity so velocity is the derivative of position, like yuo said that makes c'(t)=v(t)=?

OpenStudy (anonymous):

....how does the speed of my original question = sqrt( 1+ 64pi^2 ). i don't understand why i'm trying to find velocity right now...

OpenStudy (anonymous):

speed is the magnitude of velocity

OpenStudy (anonymous):

@TuringTest ....how does the speed of my original question = sqrt( 1+ 64pi^2 ). i don't understand why i'm trying to find velocity right now...

OpenStudy (anonymous):

NEVERMIND! got it